# R-enum Revisited: Speedup and Extension for Context-Sensitive Repeats and Net Frequencies

### v1

@@ 2.5 Burrows-Wheeler transform
The Burrows-Wheeler transform (BWT) $\mathsf{L}[1..n]$ of $T$ is the string such that $\mathsf{L}[i]=T[\mathsf{SA}[i]-1]$ for any $i\in[1..n]$. By definition, it holds that $\mathsf{lc}(x)=\{\mathsf{L}[i]\mid i\in\mathcal{I}(x)\}$ for any string $x$. The LF-mapping $\mathsf{LF}$ is the permutation on $[1..n]$ such that $\mathsf{LF}(i)$ represents the lexicographic rank of the suffix $T[\mathsf{SA}[i]-1..]$ if $\mathsf{SA}[i]\neq 1$, and otherwise $\mathsf{LF}(i)=1$. The FL-mapping $\mathsf{FL}$ is the inverse mapping of $\mathsf{LF}$. Given an integer $i$ in $\mathcal{I}(x)$ for a string $x$, we can compute $x$ by using FL-mapping $|x|$ times from $i$ because $x[k]=\mathsf{L}[\mathsf{FL}^{k}(i)]$ holds for any $k~(1\leq k\leq|x|)$, where $\mathsf{FL}^{k}(i)$ is defined recursively $\mathsf{FL}^{k}(i)=\mathsf{FL}(\mathsf{FL}^{k-1}(i))$ with the base case $\mathsf{FL}^{1}(i)=\mathsf{FL}(i)$.
Table˜2 shows an example of $\mathsf{SA}$, $\mathsf{LF}(\cdot)$, $\mathsf{FL}(\cdot)$ and $\mathsf{L}$ for $T=\mathtt{abcbbcbcabc\mathdollar}$. Figure˜1 also illustrates the suffix tree of $T=\mathtt{abcbbcbcabc\mathdollar}$ and a relationship between Weiner links and BWTs.
@@
A notable property of $\mathsf{LF}$ is that $\mathsf{LF}(i)=k_{1}+k_{2}$ holds, where $k_{1}$ is the number of characters smaller than $\mathsf{L}[i]$ in $\mathsf{L}$ and $k_{2}$ is the number of occurrences of $\mathsf{L}[i]$ in $\mathsf{L}[1..i]$. In particular, if $\mathsf{L}[p..q]$ consists of a single character, then $\mathsf{LF}(i)=\mathsf{LF}(p)+i-p$ for any $i\in[p..q]$, and $\mathsf{FL}(i^{\prime})=\mathsf{FL}(p^{\prime})+i^{\prime}-p^{\prime}$ for any $i^{\prime}\in[p^{\prime}..q^{\prime}]$ with $p^{\prime}=\mathsf{LF}(p)$ and $q^{\prime}=\mathsf{LF}(q)$. Hence, LF-mapping can be implemented in $O(r)$ space by constructing the predecessor data structure for the set of beginning positions of runs of $\mathsf{L}$ and remembering the $\mathsf{LF}(\cdot)$ values for those positions (FL-mapping can be implemented similarly). Using the predecessor data structure of Lemma˜2, we get:
Lemma 5.
@@
Nishimoto and Tabei [24] proposed a novel data structure, called the move data structure, to implement LF-mapping by a simple linear search on a list of intervals without predecessor queries. We consider a list $p_{1}..p_{2}),[p_{2}..p_{3}),\dots,[p_{\hat{r}}..p_{\hat{r}+1})$ of $\hat{r}=O(r)$ intervals such that $p_{1}=1$, $p_{\hat{r}+1}=n+1$, and for any $k\in[1..\hat{r}$], $\mathsf{L}[p_{k}..p_{k+1})$ consists of a single character and $\mathsf{LF}(p_{k})..\mathsf{LF}(p_{k+1}-1)]$ intersects with $O(1)$ intervals in the list. We let each $p_{k}..p_{k+1})$ have a pointer to the interval $p_{k^{\prime}}..p_{k^{\prime}+1})$ that contains $\mathsf{LF}(p_{k})$ and offset $o_{k}=\mathsf{LF}(p_{k})-p_{k^{\prime}}$. Then, $\mathsf{LF}(i)=\mathsf{LF}(p_{k})+i-p_{k}=p_{k^{\prime}}+o_{k}+i-p_{k}$ holds for any $i\in[p_{k}..p_{k+1})$. Given $i$ and interval $p_{k}..p_{k+1})$ that contains $i$, we can compute $\mathsf{LF}(i)$ and find the interval that contains $\mathsf{LF}(i)$ by assessing $O(1)$ intervals from $p_{k^{\prime}}..p_{k^{\prime}+1})$ to the right in the list. Note that $\mathsf{LF}(i)$ computation on the move data structure works along with the pointer to the interval in the list that contains $i$, which is handled implicitly hereafter. Under the assumption that $\sigma\leq r$ in this paper, we can construct the data structure needed for LF-mapping (and FL-mapping) in $O(r\log r)$ time [25, Appendix D]:
Lemma 6.
@@ 3 Upper bound $2r$ of net occurrences and MUSs
On the other hand, if $\mathsf{L}[i]$ is unique in $\mathsf{L}[p..q]$ and $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is unique in $T$, there is no repeat that covers $\mathsf{SA}[i]..\mathsf{SA}[i]+|x|)$, and hence, $\mathsf{SA}[i]$ is a net occurrence of $x$. ∎
Example 8.
Let us consider a repeat $\mathtt{bc}$ with $\mathcal{I}(\mathtt{bc})=[5..8]$ in the example of Table˜2. $\mathsf{SA}[6]=7$ is a net occurrence of $\mathtt{bc}$ because $\mathsf{L}[6]=\mathtt{c}$ is unique in $\mathsf{L}[5..8]=\mathtt{acab}$, and $T[\mathsf{SA}[6]..\mathsf{SA}[6]+2]=T[7..9]=\mathtt{bca}$ is unique in $T$. $\mathsf{SA}[5]=10$ and $\mathsf{SA}[7]=2$ are not net occurrences of $\mathtt{bc}$ because $\mathsf{L}[5]=\mathsf{L}[7]=\mathtt{a}$ is not a unique character in $\mathsf{L}[5..8]$. $\mathsf{SA}[7]=2$ and $\mathsf{SA}[8]=5$ are not net occurrences of $\mathtt{bc}$ because $T[2..4]=T[5..7]=\mathtt{bcb}$ is a repeat.
The following lemma was implicitly proved in [10, Theorem 24] to show the total length of all net occurrences is bounded by twice the sum of irreducible LCPs.
Lemma 9.
@@
Proof.
Let $p..q]=\mathcal{I}(x)$. If $i\in[p..q]$ is not at the run’s boundaries of $\mathsf{L}$, then at least one of $\mathsf{L}[i]=\mathsf{L}[i+1]$ with $i+1\in[p..q]$ or $\mathsf{L}[i-1]=\mathsf{L}[i]$ with $i-1\in[p..q]$ holds, which means that $\mathsf{SA}[i]$ is not a net occurrence of $x$ by Lemma˜7. $\mathsf{SA}[i]$ cannot be a net occurrence of two distinct repeats, since otherwise, the occurrence of the shorter repeat is covered by the longer one. ∎
Lemma˜9 leads to the following theorem:
Theorem 10.
@@
Proof.
Let $r_{1}$ and respectively $r_{>1}$ be the number of runs in $\mathsf{L}$ of length $1$ and longer than $1$, i.e., $r=r_{1}+r_{>1}$. Since there are $r_{1}+2r_{>1}$ positions that correspond to beginning or ending positions of runs in $\mathsf{L}$, it is clear from Lemma˜9 that the total number of net occurrences is at most $r_{1}+2r_{>1}\leq 2r$. The number is shown to be less than $2r$ by looking closer at some special positions like $1$ and $n$ on $\mathsf{L}$: Since $\mathsf{SA}[1]$ (resp. $\mathsf{SA}[n]$) cannot be a net occurrence if the first (resp. last) run of $\mathsf{L}$ is longer than $1$, only one net occurrence is charged to the first (resp. last) run of $\mathsf{L}$. ∎
A non-empty interval $[b..e]$ is called a minimal unique substring (MUS) if $T[b..e]$ is unique while $T[b+1..e]$ and T[b..e1]T[b..e-1] are repeats. In [14, 21], it has been proved that the net occurrences and MUSs are dual concepts, basically saying that for any two consecutive MUSs $[b..e]$ and $b^{\prime}..e^{\prime}]$ with $b<b^{\prime}$, $b+1$ is a net occurrence of $T[b+1..e^{\prime}-1]$. With this duality, we immediately get the following:
Corollary 11.
@@ 4.1 Rough sketch of r-enum and speedup
Conceptually, r-enum performs a breadth-first traversal on the internal nodes $\mathcal{S}^{\mathsf{rmr}}$ of the suffix tree by following Weiner links using RLBWT-based data structures. Each node $x\in\mathcal{S}^{\mathsf{rmr}}$ is visited with its suffix interval $\mathcal{I}(x)$ and length $|x|$ together with the information about its right extensions $\mathsf{rlist}(x)$, which is the list of the set $\{(c,\mathcal{I}(xc))\mid c\in\mathsf{rc}(x)\}$ sorted by the character $c$. The triplet $\mathsf{repr}(x)=(\mathcal{I}(x),\mathsf{rlist}(x),|x|)$ is called the rich representation [2, 26] of $x$, which can be stored in $O(|\mathsf{rc}(x)|)$ space.
@@
R-enum executes the above procedure in $O(r)$ space using range distinct queries on $\mathsf{L}^{\prime}[1..r]$ and RLBWT-based LF-mapping. The breadth-first traversal on $\mathcal{S}^{\mathsf{rmr}}$ can be made in $O(r)$ space because the space needed to store the rich representations for all strings in $\{ax\in\mathcal{S}\mid a\in\Sigma,x\in\mathcal{S}^{\mathsf{rmr}},|x|=t\}$ for any $t\in[0..n]$ was proved to be $O(r)$ in Section 3.3 of [26]. The total number of outputs of range distinct queries is upper bounded by $O(n)$ because each output can be charged to a distinct Weiner link. Similarly, the total number of LF-mapping needed is bounded by $O(n)$.
The original r-enum used Lemma˜5 to implement LF-mapping and runs in $O(n\log\log_{w}(n/r))$ time in total (excluding the time to output patterns for MAWs). We show that this can be improved to $O(n)$ time.
Theorem 13.
@@
Proof.
If $r=\Omega(n/\log n)$, the original r-enum with Lemma˜5 is already upper bounded by $O(n)$ because $O(n\log\log_{w}(n/r))=O(n\log\log_{w}\log n)$ and $w=\Theta(\log n)$. For the case with $r=o(n/\log n)$, we use the move data structure of Lemma˜6 to implement LF-mapping, which can be constructed in $O(r\log r)=O(n)$ time. Since LF-mapping can be done in $O(1)$ time on the move data structure, the r-enum runs in $O(n)$ time. ∎
@@ 4.2 Extension for NSMRs, SMRs, and context diversities
In this subsection, we extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR.
@@ 5 An $O(r)$-size data structure for NF-queries
In this section, we tackle the problem called Single-NF [10] on RLBWTs: We consider preprocessing the RLBWT of $T$ to construct a data structure of size $O(r)$ so that we can support NF-queries of computing $\mathsf{NF}(P)$ for any query pattern $P$.
@@
Proof.
We execute r-enum of Theorem˜14, which actually performs a breadth-first traversal of the reversed trie for $\mathcal{S}^{\mathsf{rmr}}$ in $O(n)$ time and $O(r)$ space. Our idea is to build the compacted reversed trie $\mathcal{T}$ for $\mathcal{S}^{\mathsf{nsmr}}$ during the traversal, which can be stored in $O(|\mathcal{S}^{\mathsf{nsmr}}|)=O(r)$ space because $|\mathcal{S}^{\mathsf{nsmr}}|\leq 2r$ by Lemma˜9. For each node $x$ of $\mathcal{T}$, which corresponds to a string in $\mathcal{S}^{\mathsf{nsmr}}$ or a branching node, we store the string length $|x|$ and an integer $i_{x}$ in $\mathcal{I}(x)$ so that the edge label between $x$ and its parent $y$ can be retrieved by using FL-mapping |x||y||x|-|y| times from $i_{x}$. For a node $x\in\mathcal{S}^{\mathsf{nsmr}}$, we also store the net frequency of $x$. See Figure˜2 for an illustration of $\mathcal{T}$.
@@
In total, these tasks can be done in $O(n)$ time and $O(r)$ space along with r-enum of Theorem˜14. For each branching node of $\mathcal{T}$ with $k\leq\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|)$ outgoing edges, we build a dictionary of Lemma˜1 in $O(k)$ space and $O(k)$ time w.h.p. or deterministic $O(k\log\log\min(\sigma,k))$ time so that we can decide which child to proceed by the first characters of edges in $O(1)$ time. The construction time of the dictionaries for all branching nodes is bounded by $O(|\mathcal{S}^{\mathsf{nsmr}}|)$ time w.h.p. or deterministic $O(|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time.
Given a query pattern $P$, we read the characters of $P$ from right to left and traverse $\mathcal{T}$ from the root in $O(|P|)$ time. If we can reach a node $P\in\mathcal{S}^{\mathsf{nsmr}}$, we output $\mathsf{NF}(P)$ stored in the node. ∎
We can also support queries of computing the net occurrences of a query pattern $P$ in $O(|P|+|\mathsf{NOcc}(P)|)$ time by storing $\mathsf{NOcc}(x)$ for each node $x\in\mathcal{S}^{\mathsf{nsmr}}$ in $\mathcal{T}$, which can be stored in $O(r)$ space in total due to Lemma˜9.

### v2

@@ 2.5 Burrows-Wheeler transform
The Burrows-Wheeler transform (BWT) $\mathsf{L}[1..n]$ of $T$ is the string such that $\mathsf{L}[i]=T[\mathsf{SA}[i]-1]$ for any $i\in[1..n]$. Note that $T[\mathsf{SA}[i]-1]=T[0]=\mathdollar$ for $i$ with $\mathsf{SA}[i]=1$ due to the sentinel at $T[0]$. By definition, it holds that $\mathsf{lc}(x)=\{\mathsf{L}[i]\mid i\in\mathcal{I}(x)\}$ for any string $x$. The LF-mapping $\mathsf{LF}$ is the permutation on $[1..n]$ such that $\mathsf{LF}(i)$ represents the lexicographic rank of the suffix $T[\mathsf{SA}[i]-1..]$ if $\mathsf{SA}[i]\neq 1$, and otherwise $\mathsf{LF}(i)=1$. The FL-mapping $\mathsf{FL}$ is the inverse mapping of $\mathsf{LF}$. Given an integer $i$ in $\mathcal{I}(x)$ for a string $x$, we can compute $x$ by using FL-mapping $|x|$ times from $i$ because $x[k]=\mathsf{L}[\mathsf{FL}^{k}(i)]$ holds for any $k\penalty 10000\ (1\leq k\leq|x|)$, where $\mathsf{FL}^{k}(i)$ is defined recursively $\mathsf{FL}^{k}(i)=\mathsf{FL}(\mathsf{FL}^{k-1}(i))$ with the base case $\mathsf{FL}^{1}(i)=\mathsf{FL}(i)$.
Table˜3 shows an example of $\mathsf{SA}$, $\mathsf{LF}(\cdot)$, $\mathsf{FL}(\cdot)$ and $\mathsf{L}$ for $T=\mathtt{abcbbcbcabc\mathdollar}$. Figure˜1 also illustrates the suffix tree of $T=\mathtt{abcbbcbcabc\mathdollar}$ and a relationship between Weiner links and BWTs.
@@
A notable property of $\mathsf{LF}$ is that $\mathsf{LF}(i)=k_{1}+k_{2}$ holds, where $k_{1}$ is the number of characters smaller than $\mathsf{L}[i]$ in $\mathsf{L}$ and $k_{2}$ is the number of occurrences of $\mathsf{L}[i]$ in $\mathsf{L}[1..i]$. In particular, if $\mathsf{L}[p..q]$ consists of a single character, then $\mathsf{LF}(i)=\mathsf{LF}(p)+i-p$ for any $i\in[p..q]$, and $\mathsf{FL}(i^{\prime})=\mathsf{FL}(p^{\prime})+i^{\prime}-p^{\prime}$ for any $i^{\prime}\in[p^{\prime}..q^{\prime}]$ with $p^{\prime}=\mathsf{LF}(p)$ and $q^{\prime}=\mathsf{LF}(q)$. Hence, LF-mapping can be implemented in $O(r)$ space by constructing the predecessor data structure for the set of beginning positions of runs of $\mathsf{L}$ and remembering the $\mathsf{LF}(\cdot)$ values for those positions (FL-mapping can be implemented similarly). Using the predecessor data structure of Lemma˜2.2, we get:
Lemma 2.5.
@@
Nishimoto and Tabei [2021NishimotoT_OptimTimeQueriesOnBwt_ICALP] proposed a novel data structure, called the move data structure, to implement LF-mapping by a simple linear search on a list of intervals without predecessor queries. We consider a list $p_{1}..p_{2}),[p_{2}..p_{3}),\dots,[p_{\hat{r}}..p_{\hat{r}+1})$ of $\hat{r}=O(r)$ intervals such that $p_{1}=1$, $p_{\hat{r}+1}=n+1$, and for any $k\in[1..\hat{r}$], $\mathsf{L}[p_{k}..p_{k+1})$ consists of a single character and $\mathsf{LF}(p_{k})..\mathsf{LF}(p_{k+1}-1)]$ intersects with $O(1)$ intervals in the list. We let each $p_{k}..p_{k+1})$ have a pointer to the interval $p_{k^{\prime}}..p_{k^{\prime}+1})$ that contains $\mathsf{LF}(p_{k})$ and offset $o_{k}=\mathsf{LF}(p_{k})-p_{k^{\prime}}$. Then, $\mathsf{LF}(i)=\mathsf{LF}(p_{k})+i-p_{k}=p_{k^{\prime}}+o_{k}+i-p_{k}$ holds for any $i\in[p_{k}..p_{k+1})$. Given $i$ and interval $p_{k}..p_{k+1})$ that contains $i$, we can compute $\mathsf{LF}(i)$ and find the interval that contains $\mathsf{LF}(i)$ by assessing $O(1)$ intervals from $p_{k^{\prime}}..p_{k^{\prime}+1})$ to the right in the list. Note that $\mathsf{LF}(i)$ computation on the move data structure works along with the pointer to the interval in the list that contains $i$, which is handled implicitly hereafter. Under the assumption that $\sigma\leq r$ in this paper, we can construct the data structure needed for LF-mapping (and FL-mapping) in $O(r\log r)$ time [2020NishimotoT_FasterQueriesOnBwtRuns_X3, Appendix D]:
Lemma 2.6.
@@ 3 Upper bound $2r$ of net occurrences and MUSs
On the other hand, if $\mathsf{L}[i]$ is unique in $\mathsf{L}[p..q]$ and $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is unique in $T$, there is no repeat that covers $\mathsf{SA}[i]..\mathsf{SA}[i]+|x|)$, and hence, $\mathsf{SA}[i]$ is a net occurrence of $x$.
Example 3.3.
Let us consider a repeat $\mathtt{bc}$ with $\mathcal{I}(\mathtt{bc})=[5..8]$ in the example of Table˜3. $\mathsf{SA}[6]=7$ is a net occurrence of $\mathtt{bc}$ because $\mathsf{L}[6]=\mathtt{c}$ is unique in $\mathsf{L}[5..8]=\mathtt{acab}$, and $T[\mathsf{SA}[6]..\mathsf{SA}[6]+2]=T[7..9]=\mathtt{bca}$ is unique in $T$. $\mathsf{SA}[5]=10$ and $\mathsf{SA}[7]=2$ are not net occurrences of $\mathtt{bc}$ because $\mathsf{L}[5]=\mathsf{L}[7]=\mathtt{a}$ is not a unique character in $\mathsf{L}[5..8]$. $\mathsf{SA}[7]=2$ and $\mathsf{SA}[8]=5$ are not net occurrences of $\mathtt{bc}$ because $T[2..4]=T[5..7]=\mathtt{bcb}$ is a repeat.
The following lemma was implicitly proved in [2024GuoEWZ_ExploitNewProperOfStrin_CPM, Theorem 24] to show the total length of all net occurrences is bounded by twice the sum of irreducible longest common prefixes.
Lemma 3.4.
@@
Proof 3.5.
Let $p..q]=\mathcal{I}(x)$. If $i\in[p..q]$ is not at the run’s boundaries of $\mathsf{L}$, then at least one of $\mathsf{L}[i]=\mathsf{L}[i+1]$ with $i+1\in[p..q]$ or $\mathsf{L}[i-1]=\mathsf{L}[i]$ with $i-1\in[p..q]$ holds, which means that $\mathsf{SA}[i]$ is not a net occurrence of $x$ by Lemma˜3.1. $\mathsf{SA}[i]$ cannot be a net occurrence of two distinct repeats, since otherwise, the occurrence of the shorter repeat is covered by the longer one.
Lemma˜3.4 leads to the following theorem:
Theorem 3.6.
@@
Proof 3.7.
Let $r_{1}$ and respectively $r_{>1}$ be the number of runs in $\mathsf{L}$ of length $1$ and longer than $1$, i.e., $r=r_{1}+r_{>1}$. Since there are $r_{1}+2r_{>1}$ positions that correspond to beginning or ending positions of runs in $\mathsf{L}$, it is clear from Lemma˜3.4 that the total number of net occurrences is at most $r_{1}+2r_{>1}\leq 2r$. The number is shown to be less than $2r$ by looking closer at some special positions like $1$ and $n$ on $\mathsf{L}$: Since $\mathsf{SA}[1]$ (resp. $\mathsf{SA}[n]$) cannot be a net occurrence if the first (resp. last) run of $\mathsf{L}$ is longer than $1$, only one net occurrence is charged to the first (resp. last) run of $\mathsf{L}$.
A non-empty interval $[b..e]$ is called a minimal unique substring (MUS) if $T[b..e]$ is unique while $T[b+1..e]$ and T[b..e1]T[b..e-1] are repeats. In [2011IlieS_MinimUniqueSubstAndMaxim, 2025MienoI_SpaceEfficOnlinComputOf_CPM], it has been proved that the net occurrences and MUSs are dual concepts, basically saying that for any two consecutive MUSs $[b..e]$ and $b^{\prime}..e^{\prime}]$ with $b<b^{\prime}$, $b+1$ is a net occurrence of $T[b+1..e^{\prime}-1]$. With this duality, we immediately get the following:
Corollary 3.8.
@@ 4.1 Rough sketch of r-enum and speedup
Conceptually, r-enum performs a breadth-first traversal on the internal nodes $\mathcal{S}^{\mathsf{rmr}}$ of the suffix tree by following Weiner links using RLBWT-based data structures. Each node $x\in\mathcal{S}^{\mathsf{rmr}}$ is visited with its suffix interval $\mathcal{I}(x)$ and length $|x|$ together with the information about its right extensions $\mathsf{rlist}(x)$, which is the list of the set $\{(c,\mathcal{I}(xc))\mid c\in\mathsf{rc}(x)\}$ sorted by the character $c$. The triplet $\mathsf{repr}(x)=(\mathcal{I}(x),\mathsf{rlist}(x),|x|)$ is called the rich representation [2020BelazzouguiCKM_LinearTimeStrinIndexAnd, 2021NishimotoT_REnumEnumerOfCharac_CPM] of $x$, which can be stored in $O(|\mathsf{rc}(x)|)$ space.
@@
R-enum executes the above procedure in $O(r)$ space using range distinct queries on $\mathsf{L}^{\prime}[1..r]$ and RLBWT-based LF-mapping. The breadth-first traversal on $\mathcal{S}^{\mathsf{rmr}}$ can be made in $O(r)$ space because the space needed to store the rich representations for all strings in $\{ax\in\mathcal{S}\mid a\in\Sigma,x\in\mathcal{S}^{\mathsf{rmr}},|x|=t\}$ for any $t\in[0..n]$ was proved to be $O(r)$ in Section 3.3 of [2021NishimotoT_REnumEnumerOfCharac_CPM]. The total number of outputs of range distinct queries is upper bounded by $O(n)$ because each output can be charged to a distinct Weiner link. Similarly, the total number of LF-mapping needed is bounded by $O(n)$.
The original r-enum used Lemma˜2.5 to implement LF-mapping and runs in $O(n\log\log_{w}(n/r))$ time in total (excluding the time to output patterns for MAWs). We show that this can be improved to $O(n)$ time.
Theorem 4.2.
@@
Proof 4.3.
If $r=\Omega(n/\log n)$, the original r-enum with Lemma˜2.5 is already upper bounded by $O(n)$ because $O(n\log\log_{w}(n/r))=O(n\log\log_{w}\log n)$ and $w=\Theta(\log n)$. For the case with $r=o(n/\log n)$, we use the move data structure of Lemma˜2.6 to implement LF-mapping, which can be constructed in $O(r\log r)=O(n)$ time. Since LF-mapping can be done in $O(1)$ time on the move data structure, the r-enum runs in $O(n)$ time.
@@ 4.2 Extension for NSMRs, SMRs, and context diversities
In this subsection, we extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR. Since $\mathcal{S}^{\mathsf{smr}}\subseteq\mathcal{S}^{\mathsf{nsmr}}\subseteq\mathcal{S}^{\mathsf{mr}}\subseteq\mathcal{S}^{\mathsf{rmr}}$ holds, for enumerating SMRs (resp. NSMRs), it is enough to check every element in $\mathcal{S}^{\mathsf{rmr}}$ if it is in $\mathcal{S}^{\mathsf{smr}}$ (resp. $\mathcal{S}^{\mathsf{nsmr}}$).
@@ 5 An $O(r)$-size data structure for NF-queries
In this section, we tackle the problem called Single-NF [2024GuoEWZ_ExploitNewProperOfStrin_CPM] on RLBWTs: We consider preprocessing the RLBWT of $T$ to construct a data structure of size $O(r)$ so that we can support NF-queries of computing $\mathsf{NF}(P)$ for any query pattern $P$.
@@
Proof 5.2.
We execute r-enum of Theorem˜4.4, which actually performs a breadth-first traversal of the reversed trie for $\mathcal{S}^{\mathsf{rmr}}$ in $O(n)$ time and $O(r)$ space. Our idea is to build the compacted reversed trie $\mathcal{T}$ for $\mathcal{S}^{\mathsf{nsmr}}$ during the traversal, which can be stored in $O(|\mathcal{S}^{\mathsf{nsmr}}|)=O(r)$ space because $|\mathcal{S}^{\mathsf{nsmr}}|\leq 2r$ by Lemma˜3.4. For each node $x$ of $\mathcal{T}$, which corresponds to a string in $\mathcal{S}^{\mathsf{nsmr}}$ or a branching node, we store the string length $|x|$ and an integer $i_{x}$ in $\mathcal{I}(x)$ so that the edge label between $x$ and its parent $y$ can be retrieved by using FL-mapping |x||y||x|-|y| times from $i_{x}$. For a node $x\in\mathcal{S}^{\mathsf{nsmr}}$, we also store the net frequency of $x$. See Figure˜2 for an illustration of $\mathcal{T}$.
@@
In total, these tasks can be done in $O(n)$ time and $O(r)$ space along with r-enum of Theorem˜4.4. For each branching node of $\mathcal{T}$ with $k\leq\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|)$ outgoing edges, we build a dictionary of Lemma˜2.1 in $O(k)$ space and $O(k)$ time w.h.p. or deterministic $O(k\log\log\min(\sigma,k))$ time so that we can decide which child to proceed by the first characters of edges in $O(1)$ time. The construction time of the dictionaries for all branching nodes is bounded by $O(|\mathcal{S}^{\mathsf{nsmr}}|)$ time w.h.p. or deterministic $O(|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time.
Given a query pattern $P$, we read the characters of $P$ from right to left and traverse $\mathcal{T}$ from the root in $O(|P|)$ time. If we can reach a node $P\in\mathcal{S}^{\mathsf{nsmr}}$, we output $\mathsf{NF}(P)$ stored in the node.
We can also support queries of computing the net occurrences of a query pattern $P$ in $O(|P|+|\mathsf{NOcc}(P)|)$ time by storing $\mathsf{NOcc}(x)$ for each node $x\in\mathcal{S}^{\mathsf{nsmr}}$ in $\mathcal{T}$, which can be stored in $O(r)$ space in total due to Lemma˜3.4.